# A Weighted Loss Approach to Robust Federated Learning under Data Heterogeneity

### v2

@@ Abstract
Federated learning (FL) is a machine learning paradigm that enables multiple data holders to collaboratively train a machine learning model without sharing their training data with external parties. In this paradigm, workers locally update a model and share with a central server their updated gradients (or model parameters). While FL seems appealing from a privacy perspective, it opens a number of threats from a security perspective as (Byzantine) participants can contribute poisonous gradients (or model parameters) harming model convergence. Byzantine-resilient FL addresses this issue by ensuring that the training proceeds as if Byzantine participants were absent. Towards this purpose, common strategies ignore outlier gradients during model aggregation, assuming that Byzantine gradients deviate more from honest gradients than honest gradients do from each other. However, in heterogeneous settings, honest gradients may differ significantly, making it difficult to distinguish honest outliers from Byzantine ones. In this paper, we introduce the Worker Label Alignement Loss (WoLA), a weighted loss that aligns honest worker gradients despite data heterogeneity, which facilitates the identification of Byzantines’ gradients. This approach significantly outperforms state-of-the-art methods in heterogeneous settings. In this paper, we provide both theoretical insights and empirical evidence of its effectiveness.
1 Introduction
@@
Federated learning (FL) [19, 18, 41] enables a group of machines (workers) to collaboratively train a model without sharing their private data. In this paradigm, a central server coordinates the process by sending the current model to all workers. Each worker then performs a training step on its local data and sends an update back to the server. Typically, this update is the gradient computed from a mini-batch [19, 3]. The server aggregates the gradients, updates the model, and sends it back to the workers, repeating the cycle.
@@
In an ideal setup, the server aggregates gradients by simply averaging them. However, some workers may fail or behave maliciously by sending corrupted updates (also referred to as Byzantine behaviour). Robust training seeks to achieve model performance comparable to that achieved in the absence of Byzantine workers. To this end, a common strategy is to replace simple averaging with robust aggregation methods on the server side [19, 18]. These methods are designed to reduce the impact of outliers, which are expected to include Byzantine updates. However, their effectiveness strongly depends on the assumption that honest workers hold homogeneous data distributions [18, 3, 33, 4]. In practice, this assumption may not hold [3, 24, 29, 15], as each worker typically has access to only a small, potentially unrepresentative portion of the dataset.
With heterogeneous data, the divergence among honest gradients increases, making it easier for Byzantine machines to mask their updates [3, 19, 24]. In this setting, a robust aggregator alone is not enough to resist Byzantine attacks. A common approach is to add a pre-aggregation step, such as Bucketing [24] or Nearest Neighbor Mean (NNM) [24, 3, 19, 18]. While effective under weak adversarial conditions (i.e., when the proportion of Byzantine updates is low), we show that the performance of these solutions drops significantly in stronger adversarial scenarios.
@@
We present a novel solution for robust FL in heterogeneous and highly adversarial settings. Specifically, we address label distribution skew [29, 35, 46]—a form of data heterogeneity where label distributions vary across workers, while the feature distribution for each class remains consistent across them. This scenario is common in practice [29, 35, 46]. For instance, in a mobile image recognition app, users of different ages may capture different subjects, resulting in variations in label distributions across devices. In another case, hospitals may specialize in certain types of diseases, leading to an overrepresentation of specific labels in their local patient records.
Label skewness leads to inconsistent training across workers, as each worker optimizes the model based on its own label distribution [24, 46, 4]. Local training tends to overemphasize majority classes while neglecting minority ones [46, 4]. Since these class imbalances vary across workers, local objectives become misaligned, leading to highly inconsistent updates.
@@
Rather than attempting to reduce variance of updates through pre-aggregation, we address the root cause: misaligned worker objectives. We introduce the Worker Label Alignment (WoLA) loss, a weighted loss function that aligns local training objectives. WoLA reweights point-wise errors for each worker to generate gradients as if all local datasets were drawn from the same label distribution. By steering honest gradients in a common direction, WoLA reduces variance and improves Byzantine resilience.
WoLA is model-agnostic, requires no hyperparameter tuning, no server-side dataset, no knowledge of the number of Byzantine machines, and adds no computational cost. It is also straightforward to implement and can be combined with other robustness strategies.
The only requirement is that a common training objective—represented by a probability distribution—is fixed and shared by the server once before training begins. For instance, this objective could be the global label distribution (i.e., the distribution obtained by merging all local datasets), derived either from prior knowledge or from local label distributions shared by the workers with the server. Using this information, the server estimates and broadcasts the global distribution. Despite the fact that Byzantine nodes can lie in their data distributions ahead of training, we show in this paper that their influence remains limited and that this approach still brings a significant gain in robustness compared to state-of-the-art approaches.
Our contributions can be summarized as follows:
@@
  • We introduce WoLA, a weighted loss function designed to address label skewness in highly adversarial settings (Section 5).
@@
  • We provide theoretical insights for the effectiveness of WoLA, with all proofs included in the appendix (Section 6)
  • We empirically show that WoLA significantly outperforms state-of-the-art methods through extensive experiments involving various models, datasets, aggregators, and Byzantine attacks. Our code is publicly available [^1] https://anonymous.4open.science/r/WoLA-E4CB to ensure reproducibility (Section 8).
The paper is organized as follows. Section 2 reviews the background and related work. Section 4 introduces the motivation behind the WoLA. Section 5 formally defines the loss and explores its properties. Section 6 presents theoretical justifications for its design. Section 7 describes the experimental setup, and Section 8 presents the experimental results. Finally, Section 9 concludes the paper and outlines directions for future work.
2 Background
@@
This section outlines key FL concepts and defines our threat model. All notations are summarized in Table 1.
Notation Meaning
$[k]$ Set of integers from $1$ to $k$, with $k\in\mathbb{N}$
##\## Cardinality of a set
$\mathbbm{1}$ Indicator function. For example, $\mathbbm{1}(x\leq 0)=1$ if $x\leq 0$, and 00 otherwise.
$\nabla$ Gradient
[…]
$\mathbb{E}[X]$ Expected value of the random variable $X$
$\mathbb{E}[X|A]$ Conditional expectation of the random variable $X$ given the event $A$
i.i.d. Independent and identically distributed
a.s. Almost surely is a probabilistic concept meaning that an event happens with probability $1$.
[…]
Notation Meaning
$\Phi$ Model
$n$ Number of workers
$\mathcal{H}$ Set of honest workers, with $\mathcal{H}\subset[n]$
$f$ Number of Byzantine workers, with $f=n-\#\mathcal{H}$
Notation Meaning
$C$ Number of classes
$\mathcal{D}$ Global training set, composed of pairs $(x,y)\in\mathcal{X}\times[C]$, where $x$ is an observation, and $y$ its corresponding label.
$N$ Size of the global dataset: $N=\#\mathcal{D}$
$N_{i}$ Size of worker $i$’s local dataset: $N_{i}=\#\mathcal{D}_{i}$
Notation Meaning
$\mathcal{L}$ Loss
$\ell(y,\Phi(x))$ Pointwise loss of prediction $\Phi(x)$ and label $y$
$\mathcal{L}_{i}$ Loss of worker $i$: $\mathcal{L}_{i}=\sum_{(x,y)\in\mathcal{D}_{i}}\ell\Big{(}y,\Phi(x)\Big{)}/N_{i}$, with $\sum_{i\in\mathcal{H}}\mathcal{L}_{i}\ N_{i}/N=\mathcal{L}$
$\mu^{c}$ Class-wise gradient of class $c$ in $\mathcal{D}$, i.e., $\sum_{(x,y)\in\mathcal{D}}\mathbbm{1}(y=c)\nabla\ell(c,\Phi(x))/N^{c}$
$\mu^{c}_{i}$ Class-wise gradient of class $c$ in $\mathcal{D}_{i}$, i.e., $\sum_{(x,y)\in\mathcal{D}_{i}}\mathbbm{1}(y=c)\nabla\ell(c,\Phi(x))/N^{c}_{i}$
$\mathcal{WL}$ Worker Label Alignement Loss (WoLA)
$\mathcal{WL}_{i}$ WoLA of worker $i$, with $\sum_{i\in\mathcal{H}}\mathcal{WL}_{i}\ N_{i}/N=\mathcal{WL}$
$q$ Shared training objective across workers: a probability distribution $q=(q^{1},q^{2},\ldots,q^{C})$, where $q^{i}\in[0,1]$ and $\sum_{i\in[C]}q^{i}=1$.
2.1 Byzantine-free FL
@@
We consider a distributed architecture with $n$ workers and a central server, aimed at solving a classification task with $C$ classes. The workers hold local datasets $\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{n}$. Each worker $i$ holds $N_{i}$ data points $(x,y)\in\mathcal{X}\times[C]$, where $\mathcal{X}$ is the input space, $x$ is an observation, and $y$ is its label. Given a model $\Phi$, the standard local loss function (not WoLA) for worker $i$ is:
@@
where $\ell$ is a point-wise loss function, assumed to be differentiable with respect to the model parameters.
In each communication round, the server broadcasts the model parameters to all workers. Each worker $i$ is expected to compute the gradient using its own data $\mathcal{D}_{i}$ and send back the resulting gradient $\nabla\mathcal{L}_{i}$.
2.2 Threat Model
In this paper, we focus on handling Byzantine workers under the common assumption of a trusted FL server [3, 24, 4]. We consider an adversarial setting where fewer than half the workers are Byzantine, i.e., $f<n/2$. These malicious workers can send arbitrary updates to the server and have full knowledge of all local updates, following the threat model in [3, 24, 4]. The server’s objective is to identify a stationary point of the average loss over the honest workers. That is, to find a point where $\nabla\mathcal{L}=0$, with
$\mathcal{L}:=\sum_{i\in\mathcal{H}}\mathcal{L}_{i}\frac{N_{i}}{N},$
and $N=\sum_{i\in\mathcal{H}}N_{i}$. Assuming such a point exists, it can be approximated via gradient descent and robustness strategies.
2.3 Global Dataset, Local Datasets & Label Distribution Skew
To model data heterogeneity, we consider the following label skew scenario: all data points are drawn from independent and identically distributed (i.i.d) random variables, forming the global dataset $\mathcal{D}$. These points are then randomly assigned to honest workers, resulting in local datasets $\mathcal{D}_{i}$ for $i\in\mathcal{H}$, such that $\mathcal{D}=\cup_{i\in\mathcal{H}}\mathcal{D}_{i}$. Each worker receives data with a different label distribution, introducing heterogeneity. Although Byzantine workers may also have their own data, we exclude them from the global dataset as they play no role in our analysis.
The global label distribution refers to the distribution of labels in the full dataset $\mathcal{D}$, while the local label distributions refer to label distributions in a worker’s datasets $\mathcal{D}_{i}$, for $i\in\mathcal{H}$.
2.4 Gradient Dissimilarity
Karimireddy et al. [24] introduce Gradient Dissimilarity as a measure of system robustness, lower gradient dissimilarity implies stronger robustness. It quantifies how much individual gradients deviate from the average of the honest gradients:
Gradient $\text{Gradient Dissimilarity}=\frac{1}{\#\mathcal{H}}\sum_{i\in\mathcal{H}}\|\nabla\mathcal{L}_{i}-\nabla\mathcal{L}\|^{2},$
Data heterogeneity tends to increase gradient dissimilarity [19, 18, 24].
3 Related Work
@@
3.1 Robustness Strategies with Homogeneous Data
@@
To mitigate Byzantine behavior during aggregation, a common approach is to replace simple updates averaging with robust aggregation methods [34, 19]. Examples include the coordinate-wise median (CWMed) [45], geometric median (GM) [38, 34], coordinate-wise trimmed mean (CwTM) [45], Krum [7], Multi-Krum (MKrum) [20], Minimum Diameter Averaging [20], and Comparative Gradient Elimination [22]. These methods heavily rely on the assumption that honest workers have homogeneous data. Their robustness under data heterogeneity remains unclear [18, 3, 33].
@@
Besides robust aggregators, additional strategies can enhance robustness in homogeneous settings. These approaches rely on a clean dataset held by the server to evaluate and filter worker updates [44, 9, 14, 36].
3.2 Robustness Strategies Under Data Heterogeneity
We begin by defining Gradient Dissimilarity, a robustness measure commonly used in studies addressing Byzantine resilience under data heterogeneity [19, 18, 24]. Next, we review state-of-the-art approaches in Byzantine FL, followed by methods developed for the Byzantine-free setting. Finally, we highlight the advantages of our proposed method.
3.2.1 Byzantine FL
A common strategy to reduce gradient variance is to apply a pre-aggregation step, such as Bucketing [24], Nearest-Neighbor Mixing (NNM) [3], or FoundationFL [15]. These methods act before the main aggregation rule, resulting in a two-step aggregation process—for example, applying NNM before Krum. Other methods like BOBA [4] and RAGE [12] use different strategies and do not rely on pre-aggregation.
@@ 3.2.1 Byzantine FL
Bucketing [24] randomly partitions inputs into groups (buckets), averages the inputs in each bucket, and feeds these averages into the aggregation rule. This random partitioning reduces the empirical variance of honest updates in expectation (i.e., on average). However, in some iterations, it may fail to reduce heterogeneity, giving Byzantine workers a better chance to disrupt learning [18].
@@
In contrast, NNM [3] is a deterministic pre-aggregation method that replaces each gradient with the average of its $n-f$ nearest neighbors (including itself). NNM introduces additional computational cost, which prevents its use in large-scale distributed learning tasks [18].
@@
FoundationFL [15] synthesizes additional updates at each training step. The server assigns a score to each update and replicates the highest-scoring gradient $m$ times, where $m$ is a hyperparameter set by the authors to $m=n/2$. A critical limitation of FoundationFL is its underlying assumption [^2] The score is based on the distance to the coordinate-wise minimum and maximum. Given updates $g_{1},\dots,g_{n}$, define $g_{\text{max}}=\max(g_{1},\dots,g_{n})$ and $g_{\text{min}}=\min(g_{1},\dots,g_{n})$ (coordinate-wise). The score $s_{i}$ for worker $i$ is then $s_{i}=\min(\|g_{\text{max}}-g_{i}\|,\|g_{\text{min}}-g_{i}\|)$. A high score means the update is far from the extremes.: extreme values correspond to Byzantine updates. However, this assumption is not reliable under data heterogeneity, particularly against the A Little Is Enough (ALIE) attack [5].
BOBA [4] differs from standard robustness strategies. It projects client gradients onto a $C-1$ dimensional subspace using truncated Singular Value Decomposition (SVD), estimates the $C-1$ simplex from a server-side dataset, reconstructs each client’s label distribution based on their position relative to the simplex, and discards those with abnormal values. In addition to requiring server-side data, at each round, BOBA incurs extra cost from computing the truncated SVD of the gradient matrix, which scales with both the number of clients and the gradient dimension [^3] The computational complexity of BOBA is $\mathcal{O}(kCnd)$, where $k$ is the number of truncated SVDs computed per round (typically between 3 and 5 in thei experiments), $C$ is the number of classes, $n$ the number of workers, and $d$ the gradient dimension..
RAGE [12] performs a weighted average using confidence scores assigned to gradients. These scores are computed iteratively, downweighting gradients with large projections on the principal eigenvector of the empirical covariance matrix of gradients. The covariance is recomputed with updated scores until the largest eigenvalue is sufficiently small. The computational complexity of RAGE makes it impractical for large-scale optimization [24].
3.2.2 Byzantine-free FL
In a Byzantine-free setting, non-IID data remains a major obstacle to model convergence. Several techniques have been proposed to address this challenge, including knowledge distillation [^4] Knowledge distillation encourages a student model to approximate the teacher model’s output logits, enabling it to mimic the teacher’s behavior. and logit calibration.
Some methods rely on knowledge distillation and require access to an additional test set [28, 11, 10, 30]. In contrast, [47] and [39] avoid relying on proxy datasets. They use a generator trained on the server or client to enable data-free knowledge distillation. However, these approaches require significant communication and coordination, and introduce substantial computational overhead [40, 21].
Other approaches introduce personalized loss functions based on local data distributions, including DMFL [35], FedLC [46], and FedDisk [32]. DMFL and FedLC, inspired by [8], adjust logits [^5] Logits are the outputs of the final classification layer and inputs to the softmax. before applying softmax cross-entropy, increasing margins for minority classes. These methods require the tuning of a hyperparameter with no bounded range. FedDisk uses loss with sample-level weights, learned from a global and multiple local density estimators, independent of class labels.
All these propositions either require an additional tes set, introduce significant overhead [40, 21, 32], or rely on hyperparameter tuning—a particularly difficult task in our setting: (i) without a test set, local tuning is possible, but data heterogeneity may cause inconsistent values; and (ii) since these methods assume a Byzantine-free setting, no robust protocol is provided to handle malicious behavior.
3.2.3 Our proposition
@@
To the best of our knowledge, using a loss function to improve Byzantine resilience under data heterogeneity is a novel approach. WoLA is a model-agnostic method that requires no hyperparameters, no server-side dataset, and no knowledge of the number of Byzantine machines. It adds no computational cost, is easy to implement, and can be combined with other robustness strategies. Its only requirement is a shared training objective.

### v3

@@ Abstract
Federated Learning (FL) allows multiple data holders (workers) to collaboratively train a machine learning model without sharing their private data. A central server coordinates the process by distributing the global model and receiving local gradients from workers. However, faulty or malicious (Byzantine) workers can compromise training. In heterogeneous environments, even honest gradients may differ substantially, making it challenging to distinguish them from Byzantine updates. We propose Worker Label Alignment Loss (WoLA), a weighted loss function that promotes alignment among honest gradients, thereby enhancing the detection of Byzantine behavior. Our method demonstrates superior performance compared to state-of-the-art techniques in empirical evaluations and is supported by theoretical analysis.
Codehttps://github.com/anonymous14031992/WoLA
Introduction
@@
Federated learning (FL) (Guerraoui, Gupta, and Pinot 2024b, a; Wen et al. 2023) enables multiple machines (workers) to collaboratively train a model without sharing private data. A central server coordinates the training by distributing the current global model to all workers. Each worker computes an update based on its local data and returns it to the server. The server aggregates these updates, updates the global model, and repeats the process. Typically, these updates correspond to gradients computed on local minibatches (Allouah et al. 2023).
@@
Under ideal conditions, the server aggregates gradients using simple averaging. However, some workers may fail or behave maliciously, sending corrupted (Byzantine) updates. Robust training aims to achieve performance comparable to that without Byzantine failures. A common approach replaces averaging with robust aggregation methods on the server (Guerraoui, Gupta, and Pinot 2024b, a). These methods mitigate the impact of outliers, which are expected to include Byzantine updates. Their effectiveness, however, depends heavily on honest workers having similar data distributions (Guerraoui, Gupta, and Pinot 2024a; Allouah et al. 2023; Peng, Li, and Ling 2024; Bao, Wu, and He 2024). In practice, this assumption often fails (Allouah et al. 2023; Karimireddy, He, and Jaggi 2020; Li et al. 2022; Fang et al. 2025), as each worker typically accesses only a small, potentially unrepresentative subset of the data.
In heterogeneous settings, variance among honest gradients increases, enabling Byzantine workers to better conceal their malicious updates (Allouah et al. 2023; Guerraoui, Gupta, and Pinot 2024b; Karimireddy, He, and Jaggi 2020). To enhance robustness, many methods introduce a pre-aggregation step (Guerraoui, Gupta, and Pinot 2024b, a), such as Bucketing (Karimireddy, He, and Jaggi 2020) or Nearest Neighbor Mean (NNM) (Allouah et al. 2023). While effective when the proportion of Byzantine updates is low, we show that their performance deteriorates significantly in stronger adversarial scenarios.
@@
We propose a novel approach for robust FL in heterogeneous and highly adversarial environments. Specifically, we focus on label distribution skew (Li et al. 2022; Ran, Ge, and Zhong 2021; Zhang et al. 2022)—a common form of data heterogeneity where label distributions vary across workers, but feature distributions within each class remain consistent. For example, in a mobile image recognition app, users of different ages may capture different subjects, leading to variation in label distributions across devices (Li et al. 2022; Ran, Ge, and Zhong 2021; Zhang et al. 2022). Label skew causes inconsistent training, as each worker optimizes the model according to its local label distribution (Karimireddy, He, and Jaggi 2020; Zhang et al. 2022; Bao, Wu, and He 2024), producing highly divergent updates.
@@
Instead of reducing update variance via pre-aggregation, we tackle the root cause: misaligned worker objectives. We introduce the Worker Label Alignment (WoLA) loss, a weighted loss function that aligns local training objectives across workers. WoLA reweights point-wise errors to generate gradients as if all local datasets shared the same label distribution. By steering honest gradients toward a common direction, WoLA reduces variance and improves resilience against Byzantine attacks.
Our contributions are as follows:
@@
  • We introduce WoLA, a weighted loss function designed to address label skew in highly adversarial FL settings.
@@
  • We provide theoretical analysis supporting WoLA’s effectiveness, with proofs in the appendix.
  • We empirically demonstrate that WoLA significantly outperforms state-of-the-art methods across multiple models, datasets, aggregators, and Byzantine attack scenarios.
Background
@@
This section outlines key FL concepts and defines our threat model. All notations are introduced at first use and summarized in Appendix Table 2.
Byzantine-free FL
@@
We consider a distributed FL setup with $n$ workers and a central server, tasked with a classification problem involving $C$ classes. Each worker $i$ holds a local dataset $\mathcal{D}_{i}$ of size $N_{i}$, consisting of samples $(x,y)\in\mathcal{X}\times[C]$, where $\mathcal{X}$ is the input space, $x$ an observation, and $y$ its label. Given a model $\Phi$, the local loss function for worker $i$ is:
@@
where $\ell$ is a point-wise loss, assumed differentiable with respect to model parameters.
In each communication round, the server broadcasts the current model parameters to all workers. Each worker $i$ computes the gradient $\nabla\mathcal{L}_{i}$ on its local data $\mathcal{D}_{i}$ and sends this update back to the server.
Threat Model
We assume a Byzantine setting with a trusted server and fewer than half of the workers being malicious ($f<n/2$), following the threat model in (Allouah et al. 2023; Karimireddy, He, and Jaggi 2020; Bao, Wu, and He 2024). Byzantine workers can send arbitrary updates and have full knowledge of all honest updates.
The server’s goal is to find a stationary point of the average loss over honest workers, i.e.,
$\nabla\mathcal{L}=0,\quad\text{with }\mathcal{L}:=\sum_{i\in\mathcal{H}}\frac{N_{i}}{N}\mathcal{L}_{i},\quad N=\sum_{i\in\mathcal{H}}N_{i},$
where $\mathcal{H}$ denotes the set of honest workers. Assuming such a stationary point exists, it can be approximated via gradient descent combined with robustness mechanisms.
Global Dataset, Local Datasets & Data Heterogeneity
We focus on a label skew scenario: all data points are drawn from independent and identically distributed (i.i.d) random variables, forming the global dataset $\mathcal{D}$. These points are randomly assigned to honest workers, producing local datasets $\mathcal{D}_{i}$ for $i\in\mathcal{H}$, such that $\mathcal{D}=\cup_{i\in\mathcal{H}}\mathcal{D}_{i}$. Each worker receives data with a different label distribution, introducing data heterogeneity. Although Byzantine workers may also have their own data, we exclude them from the global dataset as they play no role in our analysis.
We denote the global label distribution by $p$, corresponding to $\mathcal{D}$, and the local label distribution for worker $i$ by $p_{i}$, corresponding to $\mathcal{D}_{i}$.
Gradient Dissimilarity
Data heterogeneity increases divergence among honest gradients (Guerraoui, Gupta, and Pinot 2024b, a; Karimireddy, He, and Jaggi 2020). This divergence can be measured by Gradient Dissimilarity (Karimireddy, He, and Jaggi 2020):
$\text{Gradient Dissimilarity}=\frac{1}{\#\mathcal{H}}\sum_{i\in\mathcal{H}}\|\nabla\mathcal{L}_{i}-\nabla\mathcal{L}\|^{2},$
where $\mathcal{H}$ is the number of honest workers.
Gradient Dissimilarity is a proxy for system robustness. Lower values indicate higher robustness.
Related Work
@@
Robustness Strategies with Homogeneous Data
@@
Robust aggregation methods mitigate the impact of Byzantine failures during training (Pillutla, Kakade, and Harchaoui 2022; Guerraoui, Gupta, and Pinot 2024b). Examples include coordinate-wise median (CWMed)(Yin et al. 2018), geometric median (GM)(Small 1990; Pillutla, Kakade, and Harchaoui 2022), coordinate-wise trimmed mean (CwTM)(Yin et al. 2018), Krum(Blanchard et al. 2017), Multi-Krum (MKrum)(Guerraoui, Rouault et al. 2018), Minimum Diameter Averaging(Guerraoui, Rouault et al. 2018), and Comparative Gradient Elimination (Gupta, Liu, and Vaidya 2021). These methods strongly rely on the assumption that honest workers have homogeneous data. Their robustness under data heterogeneity remains unclear (Guerraoui, Gupta, and Pinot 2024a; Allouah et al. 2023; Peng, Li, and Ling 2024).
@@
Beyond robust aggregators, additional strategies enhance robustness in homogeneous settings by leveraging a clean dataset held by the server to evaluate and filter worker updates (Xie, Koyejo, and Gupta 2019; Cao et al. 2020; Fang et al. 2020; Regatti, Chen, and Gupta 2022).
Robustness Strategies Under Data Heterogeneity
To reduce gradient variance caused by data heterogeneity, a common approach applies a pre-aggregation step before the main aggregation rule, leading to a two-step process. Examples include Bucketing (Karimireddy, He, and Jaggi 2020), Nearest-Neighbor Mixing (NNM)(Allouah et al. 2023), and FoundationFL(Fang et al. 2025).
@@
Bucketing (Karimireddy, He, and Jaggi 2020) randomly partitions inputs into groups (buckets), averages the inputs within each bucket, and then feeds these averages into the aggregation rule. This random partitioning reduces the variance of honest updates in expectation. However, in some iterations, it may fail to reduce heterogeneity, giving Byzantine workers greater opportunity to disrupt training (Guerraoui, Gupta, and Pinot 2024a).
@@
In contrast, NNM (Allouah et al. 2023) is a deterministic pre-aggregation method that replaces each gradient with the average of its nfn-f nearest neighbors. NNM achieves optimal probabilistic robustness guarantees but has a high computational cost, limiting its scalability in large distributed systems (Guerraoui, Gupta, and Pinot 2024a).
@@
FoundationFL (Fang et al. 2025) synthesizes additional updates at each training step. The server scores each update based on its distance to the coordinate-wise extremes and replicates the highest-scoring update approximately $n/2$ times [^1] Given updates $g_{1},\dots,g_{n}$, define $g_{\text{max}}=\max(g_{1},\dots,g_{n})$ and $g_{\text{min}}=\min(g_{1},\dots,g_{n})$ (coordinate-wise). The score $s_{i}$ for worker $i$ is then $s_{i}=\min(\|g_{\text{max}}-g_{i}\|,\|g_{\text{min}}-g_{i}\|)$. A high score means the update is far from the extremes.. This scoring assumes that extreme values are Byzantine but becomes unreliable when data is heterogeneous (Baruch, Baruch, and Goldberg 2019).
Other methods, such as BOBA (Bao, Wu, and He 2024) and RAGE (Data and Diggavi 2021), adopt different strategies without relying on pre-aggregation. However, they require access to a server-side dataset and are computationally expensive, limiting their practicality in large-scale optimization (Karimireddy, He, and Jaggi 2020) [^2] At each round, BOBA requires computing several times the truncated SVD of the gradient matrix, which scales with both the number of clients and the gradient dimension..
Our proposition
@@
To the best of our knowledge, employing a loss function to improve robustness under data heterogeneity is novel. WoLA is model-agnostic, requires no hyperparameter tuning, no server-side dataset, no knowledge of the number of Byzantine workers, and adds no computational overhead. It is straightforward to implement and can be combined with other robustness strategies. WoLA only requires a common training objective shared by the server once before training begins.